OpenStudy (anonymous):

Help ! Part 1: Write an expression for the following real-world situation. Katie is moving to a new house. Small boxes hold 15 pounds each and medium boxes hold 25 pounds each. Write an expression for the total weight in the moving van if Katie uses "s" small boxes and "m" medium boxes. Part 2: Write a real-world situation of your own and a VARIABLE expression to go with it

OpenStudy (anonymous):

Small boxes hold 15 lbs each Medium boxes hold 25 lbs each Small boxes = s Medium boxes = m 15 * s = 15s 25 * m = 25m 15s + 25m = total weight.

OpenStudy (anonymous):

Part 2: Write a real-world situation of your own and a VARIABLE expression to go with it. You can use a "per pound" situation like above. ex: Amy went to the store. Pears are 5 dollars per pound. Apples are 10 dollars per pound. Write an expression to solve the total weight. p = pears and a = apples. 5p + 10a = total pounds. ^Just use something else ;)

OpenStudy (anonymous):

Ah Thank You So Much <3!

OpenStudy (anonymous):

Let me know if yu need any more!

OpenStudy (anonymous):

Okay c;

OpenStudy (anonymous):

I have a question for you.

OpenStudy (anonymous):

What Is It? :o

OpenStudy (anonymous):

Is your Dad an astronaut? Because someone took the stars from the sky and put them in your eyes.

OpenStudy (anonymous):

Aww /.\ Lol Thats So Cutee

OpenStudy (anonymous):

You are cuter! *Kisses you on the cheek*

OpenStudy (anonymous):

If you stood in front of a mirror and held up 11 roses, you would see 12 of the most beautiful things in the world.

OpenStudy (anonymous):

You're like a dictionary, you add meaning to my life!

OpenStudy (anonymous):

/.\